### v1

# HyResPINNs: Adaptive Hybrid Residual Networks for Learning Optimal Combinations of Neural and RBF Components for Physics-Informed Modeling

Abstract
Physics-informed neural networks (PINNs) are an increasingly popular class of techniques for the numerical solution of partial differential equations (PDEs), where neural networks are trained using loss functions regularized by relevant PDE terms to enforce physical constraints. We present a new class of PINNs called HyResPINNs, which augment traditional PINNs with adaptive hybrid residual blocks that combine the outputs of a standard neural network and a radial basis function (RBF) network. A key feature of our method is the inclusion of adaptive combination parameters within each residual block, which dynamically learn to weigh the contributions of the neural network and RBF network outputs. Additionally, adaptive connections between residual blocks allow for flexible information flow throughout the network. We show that HyResPINNs are more robust to training point locations and neural network architectures than traditional PINNs. Moreover, HyResPINNs offer orders of magnitude greater accuracy than competing methods on certain problems, with only modest increases in training costs. We demonstrate the strengths of our approach on challenging PDEs, including the Allen-Cahn equation and the Darcy-Flow equation. Our results suggest that HyResPINNs effectively bridge the gap between traditional numerical methods and modern machine learning-based solvers.
1 Introduction
Partial differential equations (PDEs) model a variety of phenomena across science and engineering and are traditionally solved using numerical methods such as finite difference methods (LeVeque, 2007) and finite elements (Strang et al., 1974). Physics-informed neural networks (PINNs) are meshless methods (Raissi et al., 2019; Raissi, 2018) that solve PDEs by training deep feedforward neural networks (DNNs) using PDEs as soft constraints. This traditional PINN training method poses challenges due to complicated loss landscapes arising from the PDE-based soft constraints (Krishnapriyan et al., 2021). Recent approaches to ameliorating issues include curriculum learning (Krishnapriyan et al., 2021), novel optimization techniques (Cyr et al., 2020), domain decomposition (X-PINNs) (Jagtap & Karniadakis, 2020), gradient-enhanced training (G-PINNs) (Yu et al., 2022a), or discretely-trained PINNs using RBF-FD approximations in place of automatic differentiation (DT-PINNs) (Sharma & Shankar, 2022). Further significant strides have been made in developing new DNN architectures to enhance PINNs’ representational capacity, including adaptive activation functions Jagtap et al. (2020b; a), positional embeddings Liu et al. (2020); Wang et al. (2021b), and innovative architectures Wang et al. (2021a); Sitzmann et al. (2020); Gao et al. (2021); Fathony et al. (2020); Moseley et al. (2023); Kang et al. (2023).
In addition to architectural choices, certain training pathologies exist including, spectral bias (Rahaman et al., 2019; Wang et al., 2021b), unbalanced back-propagated gradients (Wang et al., 2021a; 2022b), and causality violations (Wang et al., 2022a; 2024b). Efforts focused on improving PINNs’ training performance include loss re-weighting schemes (Wang et al., 2021a; 2022b; McClenny & Braga-Neto, 2020; 2023; Maddu et al., 2022) and adaptive resampling of collocation points, such as importance sampling (Nabian et al., 2021), evolutionary sampling (Daw et al., ), and residual-based adaptive sampling (Wu et al., 2023).
Combining traditional methods with DNNs is a new line of research which focuses on leveraging the advantages of both. For example, incorporating a Fourier feature layer to preprocess DNN inputs enhances their capacity to capture high-frequency functions (Tancik et al., 2020), while also reducing eigenvector bias in PINNs (Wang et al., 2021b; Raynaud et al., 2022). Fourier neural operators (Li et al., 2020) utilize Fourier layers in the network architecture and have become a popular approach for inverse problems. Chrysos et al. (2022; 2020) augment polynomial neural networks with DNNs, with a direct focus on classification and discriminative problems. Further explorations encompass alternative objective functions, such as those employing numerical differentiation techniques Huang & Alkhalifah (2024) and variational formulations inspired by Finite Element Methods (FEM) Kharazmi et al. (2019); Berrone et al. (2022); Patel et al. (2022), along with additional regularization terms to expedite PINNs’ convergence Yu et al. (2022b); Son et al. (2021).
Recent work has investigated deep architectures for solving PDEs through methods such as stacked DNNs (Howard et al., 2023) and residual blocks with adaptive gating parameters (Wang et al., 2024a). Wang et al. (2024a) incorporate residual blocks with adaptive skip connections, which dynamically balance the input and learned residuals at each layer. Specifically, they incorporate adaptive residual blocks which dynamically balance the contributions of both the input and residual function at each layer, while the stacked approach of  Howard et al. (2023) utilizes transfer learning to improves the ability of PINNs to handle complex, multi-scale PDEs.
While these deep residual based approaches show much promise, the increased architectural complexity, leads to higher computational costs both in terms of memory and training time—requiring careful selection of training routines to prevent instabilities or poor convergence. Finally, while Fourier features help in capturing high-frequency components, they can still struggle with discontinuous solutions or sharp interfaces, which are challenging for neural networks to approximate due to the smoothness of typical activation functions.
To address these challenges, we propose a novel class of network architectures termed HyResPINNs, that combines a standard DNN with a Radial Basis Function (RBF) network within bi-level adaptive residual blocks, such that the relative contribution of both the DNN and RBF networks are adaptively learned during training, along with the full residual block outputs. The proposed architecture leverages the strengths of both smooth and non-smooth function approximators to model complex physical systems. The standard DNN using smooth activation functions effectively captures the continuous, global behaviors of the solution, which are common in many physical phenomena. While the RBF network then captures the localized, discontinuous or sharp features in the target solution. This representational division ensures that the model can accurately capture both smooth and non-smooth components of the solution, forming a robust network capable of modeling many types of physical systems.
Our main contributions are summarized as follows:
  • Propose a novel adaptive residual architecture: We introduce a new class of physics-informed neural networks, HyResPINN, that combine standard neural networks with Radial Basis Function (RBF) networks within bi-level adaptive residual blocks.
  • Demonstrate the superiority of the block structure: We show that our residual block architecture provides significant improvements over standard approaches in capturing both smooth and non-smooth features, leading to more accurate modeling of complex physical systems.
  • Highlight the benefits of adaptivity between residual blocks: We demonstrate that the adaptive learning of contributions between the NN and RBF networks, as well as between residual blocks, results in superior performance and stability compared to non-adaptive methods.
  • Thorough empirical evaluation: Show our method HyResPINN  outperforms standard PINNs along with state of the art methods in PINNs on an array of baseline problems, confirming the general applicability of HyResPINNs.
2 Background
2.1 Physics-informed neural network
Given the spatio-temporal domain $\Omega\subset\mathbb{R}^{d}$ defined on $[0,T]\times\Omega\subset\mathbb{R}^{1+d}$ where $\Omega$ is a bounded domain in $\mathbb{R}^{d}$ with regular enough boundary $\partial\Omega$, the general form of a parabolic PDE is,
$\mathbf{u}_{t}+{\mathcal{F}}[\mathbf{u}]=\mathbf{f}\,,$ (1)
such that ${\mathcal{F}}[\cdot]$ is a linear or nonlinear differential operator, and $\mathbf{u}(t,\mathbf{x})$ denotes a unknown solution.
The general initial and boundary conditions can be then formulated as:
$\mathbf{u}(0,\mathbf{x})=\mathbf{g}(\mathbf{x}),\quad\mathbf{x}\in\Omega\,,$ (2)
$\mathcal{B}[\mathbf{u}]=0,\quad t\in[0,T],\ \mathbf{x}\in\partial\Omega\,.$ (3)
Here, $\mathbf{f}$ and $\mathbf{g}(\mathbf{x})$ are given functions with certain regularity; $\mathcal{B}[\cdot]$ denotes an abstract boundary operator, representing various boundary conditions such as Dirichlet, Neumann, Robin, and periodic conditions.
We aim at approximating the unknown solution $\mathbf{u}(t,\mathbf{x})$ by a deep neural network $\mathbf{u}_{\mathbf{\theta}}(t,\mathbf{x})$, where $\mathbf{\theta}$ denotes the set of all trainable parameters of the network (e.g., weights and biases). If a smooth activation function is employed, $\mathbf{u}_{\mathbf{\theta}}$ provides a smooth representation that can be queried for any $(t,\mathbf{x})$. The PDE residuals are defined as,
$\mathcal{R}_{{\rm int}}[\mathbf{u}_{\mathbf{\theta}}](t,\mathbf{x})=\frac{\partial\mathbf{u}_{\mathbf{\theta}}}{\partial t}(t,\mathbf{x})+\mathcal{D}[\mathbf{u}_{\mathbf{\theta}}](t,\mathbf{x})-\mathbf{f}(\mathbf{x})\,,\quad(t,\mathbf{x})\in[0,T]\times\Omega\,,$ (4)
and spatial and temporal boundary residuals, respectively, by
$\mathcal{R}_{{\rm bc}}[\mathbf{u}_{\mathbf{\theta}}](t,\mathbf{x})=\mathcal{B}[\mathbf{u}_{\theta}](t,\mathbf{x})\,,\quad(t,\mathbf{x})\in[0,T]\times\partial\Omega\,,$ (5)
and
$\mathcal{R}_{{\rm ic}}[\mathbf{u}_{\mathbf{\theta}}](\mathbf{x})=\mathbf{u}_{\theta}(0,\mathbf{x})-\mathbf{g}(\mathbf{x})\,,\quad\mathbf{x}\in\Omega\,.$ (6)
Then, we train a physics-informed model by minimizing the following composite empirical loss:
$\mathcal{L}(\mathbf{\theta}):=\underbrace{\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left|\mathcal{R}_{{\rm ic}}[\mathbf{u}_{\mathbf{\theta}}](\mathbf{x}_{ic}^{i})\right|^{2}}_{\mathcal{L}_{ic}(\mathbf{\theta})}+\underbrace{\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left|\mathcal{R}_{{\rm bc}}[\mathbf{u}_{\mathbf{\theta}}](t_{bc}^{i},\mathbf{x}_{bc}^{i})\right|^{2}}_{\mathcal{L}_{bc}(\mathbf{\theta})}+\underbrace{\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left|\mathcal{R}_{\rm int}[\mathbf{u}_{\mathbf{\theta}}](t^{i}_{r},\mathbf{x}^{i}_{r})\right|^{2}}_{\mathcal{L}_{r}(\mathbf{\theta})}\,,$ (7)
which aims to enforce the neural network function $\mathbf{u}_{\theta}$ to satisfy the PDEs (1) with initial and spatial boundary conditions (2)–(3). The training data points $\{\mathbf{x}_{ic}^{i}\}_{i=1}^{N_{ic}}$, $\{t_{bc}^{i},\mathbf{x}_{bc}^{i}\}_{i=1}^{N_{bc}}$ and $\{t_{r}^{i},\mathbf{x}_{r}^{i}\}_{i=1}^{N_{r}}$ can be the vertices of a fixed mesh or points randomly sampled at each iteration of a gradient descent algorithm.
2.2 Radial Basis Function Networks
The general form of an RBF network for the training data points $\mathbf{x}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ with corresponding function values $y_{i}$ is given by,
$f({\bf x})$ $=\sum_{j=1}^{N_{c}}c_{j}\phi(||{\bf x}-{\bf x}_{j}^{c}||),$ (8)
where ${\bf x}^{c}=\{{\bf x}_{j}^{c}\}_{j=1}^{N_{c}}$ are the center points of the radial basis functions (RBFs). Here, $\phi$ represents the radial basis function, which are commonly chosen as the Gaussian, multi-quadric, or inverse multi-quadric functions, each offering different characteristics and benefits. The elements of the kernel matrix are given by:
$\mathbf{K}_{ij}$ $=\phi(||{\bf x}_{i}-{\bf x}^{c}_{j}||),$ (9)
where ${\bf x}_{i}$ represent the data points, and (9) evaluates the RBF based on the distance between the points. The kernel matrix captures the pairwise interactions between data points through the chosen RBF, and its structure is typically symmetric and positive-definite, assuming an appropriate choice of $\phi$. The choice of center points can vary depending on the application; in some cases, the center points may coincide with the data points, but they may also be selected independently of the data points to optimize approximation accuracy or computational efficiency. The RBFs determine how much influence each center has on the input, based on the distance between the input and the center.
[…]
Figure 1: Illustration of the RBF+PINN hybrid residual block with trainable strength connections between the RBF and PINN outputs.
Gradient descent is sometimes used to optimize both the centers and the unknown weights ($c_{j}$) simultaneously. These approaches are referred to as RBF networks, where the radial basis functions are the hidden layer activations, and the RBF coefficients are the trainable network parameters. The mean-squared error between the linear combination of the hidden layer outputs and the known functions values is commonly minimized.
3 Hybrid Residual PINNs (HyResPINNs)
In this section, we describe our proposed architecture—HyResPINNs. HyResPINNs  are a novel type of residual network such that each residual block incorporates an RBF kernel along with a standard DNN.
3.1 Hybrid Residual Blocks
Figure 1 shows the hybrid residual block specifics. We denote the output of the $l$-th residual block for the input ${\bf x}^{(l)}$ as $H^{(l)}({\bf x}^{(l)})$. Formally, the forward pass of each hybrid residual block in the HyResPINN  architecture is,
$H^{(l)}({\bf x}^{(l)})=\sigma\big{(}\phi(\alpha^{(l)})F_{R}^{(l)}({\bf x}^{(l)})+(1-\phi(\alpha^{(l)}))F_{N}^{(l)}({\bf x}^{(l)})\big{)},$ (10)
such that $\sigma$ is some non-linear activation function, $\alpha^{(l)}\in\mathbb{R}$ is a trainable parameter, $\phi$ represents the sigmoid function, $F_{R}^{(l)}({\bf x}^{(l)})$ is the output of the RBF network and $F_{N}^{(l)}({\bf x}^{(l)})$ is the output from a DNN or PINN within the $l$-th block.
[…]
Figure 2: Illustration of the HyResPINN  architecture using three blocks.
Adaptive Hybrid Contribution Coefficients
The trainable parameter $\alpha^{(l)}$ controls the relative contributions of the DNN and RBF components to the output—where optimizing $\alpha$ finds the best balance between these two elements. Since the sigmoid function $\phi$ constrains its output to the range $[0,1]$, the resulting output is a convex combination of the RBF and DNN components. We further incorporate adaptable residual connection parameters between each hybrid block, denoted as $\beta^{(l)}$—similar to the approaches in (Howard et al., 2023) and (Wang et al., 2024a)
However, unlike (Wang et al., 2024a), where their adaptable parameters are initialized to zero to force the network to learn non-linearities from scratch, we initialize each $\alpha^{(l)}=0.5$, ensuring equal contributions from both components at the start of training, and each $\beta^{(l)}=1$. When $\alpha=1$, only the RBF network contributes; when $\alpha=0$, only the DNN contributes. The optimal choice of $\alpha$ depends on the problem characteristics. For problems with large regions of smoothness, the model might favor a lower $\alpha$, assigning more weight to the DNN components in each residual block. Conversely, for problems involving sharp transitions or discontinuities, a higher $\alpha$ (favoring the RBF) may be preferable. In hybrid problems with smooth and non-smooth regions, $\alpha$ will likely fall between 0.4 and 0.6, providing a balanced combination of the two networks.
To encourage smoother solutions and prevent the network from introducing excessive non-linearity, we add a regularization term to each block’s trainable parameter $\alpha$. This regularization penalizes large values of $\alpha$, effectively controlling the contribution of non-linearity from the RBF components. Specifically, the total loss function is defined as:
$\mathcal{L}=\lambda_{ic}\mathcal{L}_{ic}({\boldsymbol{\theta}})+\lambda_{bc}\mathcal{L}_{bc}({\boldsymbol{\theta}})+\lambda_{r}\mathcal{L}_{r}({\boldsymbol{\theta}})+\lambda_{p}\sum_{i=1}^{N_{blocks}}\alpha_{i}^{2},$ (11)
where $\mathcal{L}_{ic}$ and $\mathcal{L}_{bc}$ represent the loss terms for the initial and boundary conditions, $\mathcal{L}_{r}$ represents the residual loss, and the final term $\lambda_{p}\sum_{i=1}^{N_{blocks}}\alpha_{i}^{2}$ acts as an $L_{2}$ regularization on the $\alpha$ parameters. The regularization helps balance the contributions of the smooth neural network and non-linear RBF components, promoting smoother solutions and more stable training.
Figure 3: 2D Allen-Cahn equation: Comparison of (a subset of) the learned RBF kernels for a standard RBF network (top row) and the proposed hybrid RBF+NN residual block approach (bottom row). Each subplot shows the RBF value corresponding to different learned RBF centers in the input domain, marked by red crosses.
Adaptive RBF Block Kernel
The output of the RBF kernel $F_{R}^{(l)}({\bf x}^{(l)})$ can be described as,
$F_{R}^{(l)}({\bf x}^{(l)})=K({\bf x})W,$ (12)
such that the kernel matrix $K$ is defined as (9) where the kernel matrix $K$ is based on the Wendland $C^{4}$ kernel in 2D. In this work, we focus on isotropic version of the Wendland kernel.
The isotropic Wendland $C^{4}$ kernel for input ${\bf x}$ and the $i$-th center is defined as:
$\phi_{i}(\mathbf{x})=\left(1-\frac{\|{\bf x}-{\bf x}^{c}_{i}\|}{\tau_{i}}\right)_{+}^{6}\left(35\left(\frac{\|{\bf x}-{\bf x}^{c}_{i}\|}{\tau_{i}}\right)^{2}+18\frac{\|{\bf x}-{\bf x}^{c}_{i}\|}{\tau_{i}}+3\right),$ (13)
where ${\bf x}^{c}_{i}$ is the center of the $i$-th RBF, $\tau$ is a scaling parameter, and $\|{\bf x}-{\bf x}^{c}_{i}\|$ represents the Euclidean distance between ${\bf x}$ and ${\bf x}^{c}_{i}$. The Wendland kernel is compactly supported, meaning $\phi_{i}(\mathbf{x})=0$ for $\|{\bf x}-{\bf x}^{c}_{i}\|\geq\tau_{i}$, which leads to sparse kernel matrices and computational efficiency. Each $\tau_{i}$ is a trainable parameter in the network and optimize through gradient descent along with all other network parameters. Figure 3 demonstrates learned scale parameters within an optimized RBF kernel.
Block Neural Network
Following the convention of Cyr et al. (2020), we represent the output of the DNN $F_{N}^{(l)}({\bf x}^{(l)})$ $F_{N}^{(l)}\in\mathbb{R}^{d}\rightarrow\mathbb{R}$ of width $w$, as a linear combination of adaptive basis functions given by
$F_{N}^{(l)}({\bf x};a,\theta^{h})=\sum_{i=1}^{w}a_{i}\sigma_{i}{\left({\bf x};\theta^{h}\right)},$ (14)
where each $a_{j}$ for $j=1,..,w$ and $\theta^{h}$ constitute the weights and biases in the last layer and hidden layers respectively, forming the set of all network parameters $\theta$. Then, each $\sigma_{j}$ are non-linear smooth activation functions such as Tanh acting on the outputs of the hidden layers. We choose to use a standard DNN architecture, but any DNN architecture such as ResNets, would work. The parameters $\theta$ are computed through some iterative optimization technique. In this work, we use variants of gradient descent methods such as ADAM Kingma & Ba (2015) and L-BFGS Liu & Nocedal (1989).
Figure 2 shows a visualization of our full model architecture with adaptive residual block skip connections, along with the input block structure used to lift the inputs to the desired higher dimension, and the output neural network block used to project each block’s output down to the output dimension.
Problem Domain Boundary Cond. PINN ResPINN Expert Stacked PirateNet HyResPINN
Allen-Cahn 1D Space/Time Periodic $0.526$ $0.0027$ $0.00386$ $0.00587$ $0.00022$ $\mathbf{9.62\times 10^{-5}}$
DarcyFlow 2D Annulus Neumann $0.00075$ $0.00046$ $0.0005$ $0.0009$ $8.71\times 10^{-5}$ $\mathbf{5.44\times 10^{-5}}$
(smooth coefficients) Dirichlet $0.0020$ $0.0014$ $0.00012$ $0.0041$ $0.00017$ $\mathbf{6.0\times 10^{-5}}$
3D Annulus Neumann $0.0022$ $0.012$ $0.021$ $0.054$ $0.039$ $\mathbf{0.0012}$
Dirichlet $0.0085$ $0.0061$ $0.0011$ $0.039$ $0.0013$ $\mathbf{0.0011}$
(rough coefficients) 2D Box Neumann $6.85\times 10^{-5}$ $2.69\times 10^{-5}$ $0.00011$ $0.00015$ $5.44\times 10^{-5}$ $\mathbf{1.05\times 10^{-5}}$
Table 1: Relative $L^{2}$ test error results for various PDE problems and baseline methods.
4 Experimental Results and Discussion
In this section, we demonstrate the effectiveness of the proposed HyResPINNs  architecture across a diverse collection of benchmark problems compared to the leading baseline methods. We show that HyResPINNs  consistently produce lower relative $L^{2}$ errors between the predicted and ground truth solutions for the same training set size when compared to the applicable baselines trained under the same experimental procedures and hyperparameters. We compare each method on the 1D non-linear hyperbolic Allen-Cahn equation and the Darcy Flow equation in two and three dimensions with both Dirichlet and Neumann boundary conditions. Our main results are summarized in Table 1.
Figure 4: 2D Allen-Cahn equation: Comparison of the predicted solutions for the Allen-Cahn equation using HyResPINN, standard PINN, and RBF network models. The top row shows the predicted solutions at for HyResPINN (left), standard PINN (center), and RBF network (right). The second row illustrates the absolute error between the predicted solutions and the exact solution. The bottom row shows the predicted solutions for three time steps comparing the exact solution (green), HyResPINN (orange), and standard PINN (dashed blue). The HyResPINN model effectively captures both the smooth and sharp features of the solution, while the standard PINN struggles with sharp transitions, leading to larger errors in these regions.
Baseline Methods. We compare the proposed HyResPINN  against a suite of baseline methods under exactly the same hyper-parameter settings in each example. Baselines include the standard PINN as originally formulated in (Raissi et al., 2019) (PINN), a PINN based on the architectural guidelines proposed in Wang et al. (2023) (ExpertPINNs), PINNs with residual connections (ResPINNs), PirateNets Wang et al. (2024a), and the stacked PINN approach of (Howard et al., 2023) (StackedPINNs). To ensure a fair comparison, we implemented the approaches detailed in (Raissi et al., 2019; Wang et al., 2024a; 2023; Howard et al., 2023) following the architectural details provided in each.
Experimental Setup. We follow similar experimental design procedures as those described in Wang et al. (2022a; 2023; 2024a). A full description of the experimental setups are included in Appendix A, and exact hyper-parameter setting are detailed in Table 2. We train all models on Nvidia A100 GPU running Centos 7.2. Code for our methods and all compared baseline approaches is written in libtorch $1.11$ (the C++ version of PyTorch) and will be made publicly available upon publication.
4.1 1D Allen-Cahn
We first focus on the Allen-Cahn equation, a challenging benchmark for conventional PINN models that has been extensively studied in recent literature  (Wight & Zhao, 2020; Wang et al., 2022a; Daw et al., ). For simplicity, we consider the one-dimensional case with a periodic boundary condition with $t\in[0,1]$, and $x\in[-1,1]$:
$u_{t}-0.0001u_{xx}+5u^{3}-5u=0\,,$
$u(0,x)=x^{2}\cos(\pi x)\,,$
$u(t,-1)=u(t,1)\,,\;u_{x}(t,-1)=u_{x}(t,1)\,.$
The Allen-Cahn PDE is an interesting benchmark for PINNs as it introduces periodic boundary conditions, and because it is a “stiff” PDE that challenges PINNs to approximate solutions with sharp space and time transitions.
Hybrid residual blocks improve prediction accuracy.
We first demonstrate that the hybrid residual block structure introduced in our method enhances the ability to capture both smooth and sharp features in the 2D Allen-Cahn solution. As shown in Figure 4, the proposed approach outperforms competing baselines, which struggle to accurately represent the sharp transitions in the solution. This difficulty arises from the smoothness constraints of standard neural network architectures. In contrast, our method effectively captures these sharp transitions by leveraging a combination of smooth neural networks and adaptive non-smooth RBF kernels, allowing the model to balance between global smooth behavior and localized discontinuities. Figure 3 shows a random subset of the learned RBF kernels captured by the proposed hybrid approach. Specifically, it captures similar RBF structures (to a standard RBF network), but benefits from the additional flexibility provided by the NN component, allowing for more nuanced function approximation in regions requiring a balance of smoothness and sharp transitions. This is likely apparent in the on average smaller smaller kernels learned by the hybrid block.
Stacking structure of residual blocks improve prediction accuracy.
Figure 5: Allen-Cahn equation: Comparison of the mean relative $L^{2}$ error using various methods as a function of the number of hidden layers.
We next show that the stacking structure employed by the HyResPINN  in each hybrid residual block structure introduced here further helps to capture the smooth and sharp features in the Allen-Cahn solution. Figure 5 shows the predicted solutions between the proposed method and the competing baselines. The competing methods need help accurately capture the sharp transitions in the solutions. In contrast, our approach more accurately captures these sharp transitions due to the conjunction of the smooth neural network architecture and the learned non-smooth RBF kernels.
Figure 6: Allen-Cahn equation: Comparison of mean relative $L^{2}$ error across various methods, plotted against the number of training iterations (left) and the mean training time (right). Vertical bars in right plot indicate the time for which each method achieved an error of $210^{-}$ 2.
We next evaluate the effect of network depth on predictive accuracy for different architectures. Figure 5 HyResPINN consistently achieves the lowest error across different network depths. The performance of other methods shows varying sensitivity to the number of hidden layers, with HyResPINN offering the most robust performance. Further, the left plot in Figure 6 shows that HyResPINN and PiratePINN achieve the lowest error over iterations, with HyResPINN consistently reducing the error more effectively than the other methods. The right plot demonstrates that—although HyResPINN has a slightly higher training cost—it significantly outperforms other methods in accuracy, particularly for long training times. These results highlight the robustness and efficiency of HyResPINN in solving complex PDEs.
Remark
While attempting to replicate the results of (Wang et al., 2024a), using the reported hyperparameters and implementation details, we observed a slight performance gap between the results obtained in our implementation and those listed in their paper. The observed discrepancy in the final model’s performance could stem from several factors such as subtle variations in the implementation environment—such as hardware differences, software library versions, or random seed initialization. Despite the differences between our replication results and those reported in (Wang et al., 2024a), the comparison between our architecture and the replicated version of their approach remains fair, as both were tested under identical conditions using the same dataset, hyperparameters, and training setup (including hardware and software environment). Our findings indicate that our architecture outperforms the replicated version of their model in many cases, along with all other replicated baseline methods.
4.2 Darcy Flow
In this section, we describe the Darcy Flow problem, an elliptic boundary value problem given by,
$-\nabla\cdot\mu\nabla\phi=f\quad\text{in }\Omega$ (15)
$\phi=u\quad\text{on }\Gamma_{D}$
${\bf n}\cdot\mu\nabla\phi=g\quad\text{on }\Gamma_{N}$
where $\Gamma_{D}$ and $\Gamma_{N}$ denote Dirichlet and Neumann parts of the boundary $\Gamma$, respectively, $\mu$ is a symmetric positive definite tensor describing a material property and $f$, $u$ and $g$ are given data.
In important applications such as porous media flow, heat transfer, and semiconductor devices, the flux ${\bf u}=-\mu\nabla\phi$ is the variable of primary interest. Specifically, during the fabrication of semiconductor devices, impurities are introduced into the silicon substrate to alter its electrical properties—a process known as doping. The diffusion of dopants can be described by equations that are similar in form to Darcy’s law, where the dopant concentration gradient drives the diffusion process. In semiconductor doping, impurities are introduced into the substrate to change its electrical properties. This involves diffusion, where the concentration gradient drives the movement of dopant atoms into the semiconductor material. While Darcy’s law governs fluid flow through porous media based on pressure gradients, the diffusion of dopants in semiconductors is guided by concentration gradients, analogous to how fluids move through porous materials in Darcy’s law.
Figure 7: 2D and 3D Smooth Darcy Flow equation: Comparison of the mean relative $L^{2}$ errors across each baseline method for the 2D Darcy Flow problem (left) and the 3D Darcy flow problem (right) plotted against the number of training collocation points. Solid lines show the solution error, while dashed lines show the x-directional flux errors.
4.2.1 Annulus Domain with Smooth Coefficients
We initially showcase the convergence rate of our method by employing smooth manufactured solutions. We consider two scenarios: one within an annular 2D domain, and another within a 3D domain formed by extruding the 2D geometry in the z-direction, resulting in a cylinder with a height of two. We present a convergence study analyzing the error rate as the number of points increases. The exact solutions are set as,
$u(x,y)=\sin(x)\sin(y),\;\;\;\;u(x,y,z)=\sin(x)\sin(y)\sin(z),$ (16)
for the 2D and 3D cases, respectively. By substituting these exact solutions into the model problem we define the source term and the boundary data. We solve the discrete problem with either Dirichlet or Neumann boundary conditions. Convergence results are shown in Figure 7 for the two and three dimensional problems and demonstrate that HyResPINNs  remain robust to boundary condition type, size of training set (achieves lowest errors compared to the baselines for fewer training points), and problem dimension.
4.2.2 2D Box Domain with Rough Coefficients
A defining feature of methods for solving the Darcy Flow equation 15, lies in their proficiency in accurately depicting the flux variable in scenarios where the coefficients $\mu$ are discontinuous. In these instances, the flux’s normal component maintains continuity across material interfaces, whereas the tangential component may exhibit discontinuities. In contrast, collocated methods like Galerkin, stabilized Galerkin, and least-squares finite elements often fail to replicate this physical behavior, typically resulting in oscillations at the interface. In this section, we demonstrate that HyResPINNs  successfully provides physically accurate flux approximations for problems characterized by discontinuous coefficients.
The initial example presented is the well-documented "five strip problem" Nakshatrala et al. (2006); Masud & Hughes (2002), which serves as a conventional manufactured solution test to evaluate a method’s capability to preserve the continuity of normal flux. The prescribed exact solution on domain $\Omega=[0,1]^{2}$ is given by,
$\phi_{ex}=1-x,\;\;\;\;\text{and}\;\;\;\Gamma_{N}=\Gamma,$ (17)
such that $\Omega$ is divided in five equal strips,
[…]
with different $\mu_{i}$ on each $\Omega_{i}$ such that $\mu_{1}=16,\;\;\mu_{2}=6,\;\;\mu_{3}=1,\;\;\mu_{4}=10,\;\;\mu_{5}=2$. We solve the discrete problem with Neumann boundary conditions. Convergence results are shown in Figure 8 and show that HyResPINNs  provide accurate solutions, albeit with larger training sizes.
Figure 8: 2D Rough Darcy Flow equation: Comparison of the mean relative $L^{2}$ errors across various methods, plotted against the number of training collocation points. Solid lines show the solution error, while dashed lines show the x-directional flux errors.
5 Conclusion
In this work, we introduced HyResPINNs, a novel class of physics-informed neural networks that incorporate adaptive hybrid residual blocks combining the strengths of standard neural networks and radial basis function (RBF) networks. Our architecture effectively captures continuous and discontinuous features by leveraging smooth and non-smooth function approximators. The adaptive combination parameters within each block allow the model to balance the contributions of neural and RBF components during training. Furthermore, including Wendland kernels enhances the model’s ability to handle sharp transitions while maintaining computational efficiency. Our experiments demonstrate that HyResPINNs  outperform traditional PINNs and state-of-the-art methods in accuracy and robustness, particularly for problems involving mixed smooth and non-smooth regions. This work increases the flexibility and generalizability of PINNs by bridging the gap between classical numerical methods and DNN-based approaches for solving PDEs.
Appendix A Experimental Design
We use mini-batch gradient descent, where collocation points are randomly sampled during training iteration, for all Allen-Cahn experiments, and we use full-batch gradient descent (but varying training set size) for all other experiments. We use the Adam optimizer Kingma & Ba (2015), and follow the learning rate schedule of Wang et al. (2023) which starts with a linear warm-up phase of $5,000$ iterations, starting from zero and gradually increasing to $10^{-3}$, followed by an exponential decay at a rate of $0.9$. Following the best practices described in Wang et al. (2023), we also employ a learning rate annealing algorithm Wang et al. (2021a; 2023) to balance losses and causal training Wang et al. (2022a; 2023) to mitigate causality violation in solving time-dependent PDEs and apply exact periodic boundary conditions Dong & Ni (2021) when applicable. We use the hyperbolic tangent activation functions and initialize each network’s parameters using the Glorot normal scheme, unless otherwise specified. We ran five random trials for each test, and report the mean values achieved in each plot and table.
Parameter Value
Architecture
Number of layers 9
Number of channels 128
Activation Tanh
Fourier feature scale 2.0
Random weight factorization $\mu=1.0,\sigma=0.1$
Learning rate schedule
Initial learning rate $10^{-3}$
Decay rate 0.9
Decay steps $5\times 10^{3}$
Warmup steps $5\times 10^{3}$
Training
Training steps $3\times 10^{5}$
Weighting
Weighting scheme Gradient Norm Wang et al. (2022b; 2023)
Causal tolerance 1.0
Number of chunks 32
Table 2: Hyper-parameter configurations for experiments.

### v2

@@
-HyResPINNs: Adaptive Hybrid Residual Networks for Learning Optimal Combinations of Neural and RBF Components for Physics-Informed Modeling
+HyResPINNs: Hybrid Residual Networks for Adaptive Neural and RBF Integration in Solving PDEs
 Abstract
-Physics-informed neural networks (PINNs) are an increasingly popular class of techniques for the numerical solution of partial differential equations (PDEs), where neural networks are trained using loss functions regularized by relevant PDE terms to enforce physical constraints.
-We present a new class of PINNs called HyResPINNs, which augment traditional PINNs with adaptive hybrid residual blocks that combine the outputs of a standard neural network and a radial basis function (RBF) network.
-A key feature of our method is the inclusion of adaptive combination parameters within each residual block, which dynamically learn to weigh the contributions of the neural network and RBF network outputs.
-Additionally, adaptive connections between residual blocks allow for flexible information flow throughout the network.
-We show that HyResPINNs are more robust to training point locations and neural network architectures than traditional PINNs.
-Moreover, HyResPINNs offer orders of magnitude greater accuracy than competing methods on certain problems, with only modest increases in training costs.
-We demonstrate the strengths of our approach on challenging PDEs, including the Allen-Cahn equation and the Darcy-Flow equation.
-Our results suggest that HyResPINNs effectively bridge the gap between traditional numerical methods and modern machine learning-based solvers.
+Physics-informed neural networks (PINNs) have emerged as a powerful approach for solving partial differential equations (PDEs) by training neural networks with loss functions that incorporate physical constraints.
+In this work, we introduce HyResPINNs, a novel class of PINNs featuring adaptive hybrid residual blocks that integrate standard neural networks and radial basis function (RBF) networks.
+A distinguishing characteristic of HyResPINNs is the use of adaptive combination parameters within each residual block, enabling dynamic weighting of the neural and RBF network contributions.
+Our empirical evaluation of a diverse set of challenging PDE problems demonstrates that HyResPINNs consistently achieve superior accuracy to baseline methods.
+These results highlight the potential of HyResPINNs to bridge the gap between classical numerical methods and modern machine learning-based solvers, paving the way for more robust and adaptive approaches to physics-informed modeling.
 1 Introduction
 Partial differential equations (PDEs) model a variety of phenomena across science and engineering and are traditionally solved using numerical methods such as finite difference methods (LeVeque, 2007) and finite elements (Strang et al., 1974).
-Physics-informed neural networks (PINNs) are meshless methods (Raissi et al., 2019; Raissi, 2018) that solve PDEs by training deep feedforward neural networks (DNNs) using PDEs as soft constraints.
+Physics-informed neural networks (PINNs) are meshless methods (Raissi et al., 2019b; Raissi, 2018) that solve PDEs by training deep feedforward neural networks (DNNs) using PDEs as soft constraints.
 This traditional PINN training method poses challenges due to complicated loss landscapes arising from the PDE-based soft constraints (Krishnapriyan et al., 2021).
@@
 (2023).
-In addition to architectural choices, certain training pathologies exist including, spectral bias (Rahaman et al., 2019; Wang et al., 2021b), unbalanced back-propagated gradients (Wang et al., 2021a; 2022b), and causality violations (Wang et al., 2022a; 2024b).
-Efforts focused on improving PINNs’ training performance include loss re-weighting schemes (Wang et al., 2021a; 2022b; McClenny & Braga-Neto, 2020; 2023; Maddu et al., 2022) and adaptive resampling of collocation points, such as importance sampling (Nabian et al., 2021), evolutionary sampling (Daw et al., ), and residual-based adaptive sampling (Wu et al., 2023).
-Combining traditional methods with DNNs is a new line of research which focuses on leveraging the advantages of both.
-For example, incorporating a Fourier feature layer to preprocess DNN inputs enhances their capacity to capture high-frequency functions (Tancik et al., 2020), while also reducing eigenvector bias in PINNs (Wang et al., 2021b; Raynaud et al., 2022).
-Fourier neural operators (Li et al., 2020) utilize Fourier layers in the network architecture and have become a popular approach for inverse problems.
-Chrysos et al.
-(2022; 2020) augment polynomial neural networks with DNNs, with a direct focus on classification and discriminative problems.
-Further explorations encompass alternative objective functions, such as those employing numerical differentiation techniques Huang & Alkhalifah (2024) and variational formulations inspired by Finite Element Methods (FEM) Kharazmi et al.
+Combining traditional methods with DNNs is a growing area of research that seeks to leverage the advantages of both approaches.
+For instance, Fourier feature layers have been incorporated into DNNs to improve their ability to represent high-frequency functions (Tancik et al., 2020) while also reducing eigenvector bias in PINNs (Wang et al., 2021b; Raynaud et al., 2022).
+Fourier neural operators (Li et al., 2020) extend this idea by incorporating Fourier layers within network architectures, which have proven useful for solving inverse problems.
+Similarly, physics-informed radial basis networks (PIRBNs) have been developed as efficient local approximators that integrate domain knowledge, particularly excelling in solving PDEs with high-frequency features or ill-posed domains (Bai et al., 2023; Fu et al., 2024).
+Other works have explored polynomial-based neural networks (Chrysos et al., 2022; 2020) and alternative optimization techniques, such as numerical differentiation methods Huang & Alkhalifah (2024) and variational formulations inspired by Finite Element Methods (FEM) Kharazmi et al.
 (2019); Berrone et al.
 (2022); Patel et al.
-(2022), along with additional regularization terms to expedite PINNs’ convergence Yu et al.
+(2022).
+Additional regularization strategies have also been proposed to improve the convergence behavior of PINNs Yu et al.
 (2022b); Son et al.
 (2021).
-Recent work has investigated deep architectures for solving PDEs through methods such as stacked DNNs (Howard et al., 2023) and residual blocks with adaptive gating parameters (Wang et al., 2024a).
-Wang et al.
-(2024a) incorporate residual blocks with adaptive skip connections, which dynamically balance the input and learned residuals at each layer.
-Specifically, they incorporate adaptive residual blocks which dynamically balance the contributions of both the input and residual function at each layer, while the stacked approach of  Howard et al.
-(2023) utilizes transfer learning to improves the ability of PINNs to handle complex, multi-scale PDEs.
-While these deep residual based approaches show much promise, the increased architectural complexity, leads to higher computational costs both in terms of memory and training time—requiring careful selection of training routines to prevent instabilities or poor convergence.
-Finally, while Fourier features help in capturing high-frequency components, they can still struggle with discontinuous solutions or sharp interfaces, which are challenging for neural networks to approximate due to the smoothness of typical activation functions.
-To address these challenges, we propose a novel class of network architectures termed HyResPINNs, that combines a standard DNN with a Radial Basis Function (RBF) network within bi-level adaptive residual blocks, such that the relative contribution of both the DNN and RBF networks are adaptively learned during training, along with the full residual block outputs.
-The proposed architecture leverages the strengths of both smooth and non-smooth function approximators to model complex physical systems.
-The standard DNN using smooth activation functions effectively captures the continuous, global behaviors of the solution, which are common in many physical phenomena.
-While the RBF network then captures the localized, discontinuous or sharp features in the target solution.
-This representational division ensures that the model can accurately capture both smooth and non-smooth components of the solution, forming a robust network capable of modeling many types of physical systems.
-Our main contributions are summarized as follows:
-  • Propose a novel adaptive residual architecture: We introduce a new class of physics-informed neural networks, HyResPINN, that combine standard neural networks with Radial Basis Function (RBF) networks within bi-level adaptive residual blocks.
-  • Demonstrate the superiority of the block structure: We show that our residual block architecture provides significant improvements over standard approaches in capturing both smooth and non-smooth features, leading to more accurate modeling of complex physical systems.
-  • Highlight the benefits of adaptivity between residual blocks: We demonstrate that the adaptive learning of contributions between the NN and RBF networks, as well as between residual blocks, results in superior performance and stability compared to non-adaptive methods.
-  • Thorough empirical evaluation: Show our method HyResPINN  outperforms standard PINNs along with state of the art methods in PINNs on an array of baseline problems, confirming the general applicability of HyResPINNs.
+Recent studies have investigated the use of deep architectures for solving PDEs through methods such as stacked DNNs (Howard et al., 2023) and residual networks with adaptive gating parameters (Wang et al., 2024) (PirateNets).
+PirateNets introduce adaptive residual blocks that dynamically balance the contributions of input features and learned residuals at each layer.
+Meanwhile, stacked PINNs utilize transfer learning to enhance PINNs’ ability to solve complex, multi-scale PDEs.
+Despite these advances, several challenges remain.
+Residual-based deep architectures often introduce significant computational costs due to increased memory and training time requirements.
+These approaches require careful tuning of training routines to prevent instabilities or slow convergence.
+Additionally, while Fourier-based methods improve the ability of DNNs to capture high-frequency components, they still struggle with discontinuous solutions or sharp interfaces—structures that neural networks inherently struggle to approximate due to the smoothness constraints imposed by typical activation functions.
+Moreover, approaches such as PirateNets rely on residual connections initialized to zero, forcing the network to learn complex function structures from scratch rather than leveraging prior knowledge from earlier layers.
+Since these methods rely entirely on DNN-based architectures, they lack explicit mechanisms for handling both smooth and non-smooth function components in a structured way.
+To address these challenges, we propose a novel class of network architectures termed HyResPINNs, which introduces a hybrid function approximation framework by integrating standard DNNs with Radial Basis Function (RBF) networks within bi-level adaptive residual blocks.
+Unlike existing residual architectures, HyResPINN adaptively learns both the relative contributions of the DNN and RBF components within each block, as well as the full residual block outputs throughout training.
+This hybrid formulation enables the model to leverage the strengths of both smooth and non-smooth function approximators to more effectively capture complex physical systems.
+The standard DNN, utilizing smooth activation functions, is well-suited for learning continuous, global behaviors commonly observed in many physical phenomena.
+Meanwhile, the RBF network component provides localized function approximation, effectively capturing sharp interfaces, discontinuities, and high-gradient regions that are difficult for purely neural network-based approaches to represent.
+By combining these two complementary representations in an adaptive manner, the proposed architecture ensures that the model can simultaneously handle both smooth and non-smooth components of the solution.
+This results in a more robust and expressive model for solving PDEs across a wide range of problem domains.
+We highlight the following key contributions:
+  • A novel hybrid residual architecture: We introduce HyResPINNs, a new class of PINNs that combine deep neural networks with radial basis function networks, enabling an adaptive balance between global smoothness and localized feature representations.
+  • Bi-level adaptivity through convex combination residual learning: Unlike prior residual-based approaches that use fixed or additive skip connections, we propose a convex combination formulation that dynamically learns the optimal contribution of each residual block output and its input, resulting in improved stability and adaptability during training.
+  • Robust handling of both smooth and non-smooth features: The proposed architecture effectively captures both globally smooth behaviors and sharp, localized structures by leveraging the complementary strengths of neural networks and RBFs, outperforming existing methods in modeling multi-scale and discontinuous solutions.
+  • Improved training stability through regularization: We introduce an explicit regularization mechanism on the adaptive gating parameters, preventing excessive dominance of either the DNN or RBF components, thereby ensuring a well-balanced hybrid function approximation.
+  • Comprehensive empirical evaluation: We demonstrate the effectiveness of HyResPINNs  across a range of benchmark PDE problems, showing superior accuracy, efficiency, and stability compared to standard PINNs and state-of-the-art baseline architectures.
+Through these contributions, we establish HyResPINNs  as a principled and effective approach for learning solutions to PDEs, providing a flexible framework that generalizes well across different problem domains.
 2 Background
-2.1 Physics-informed neural network
-Given the spatio-temporal domain $\Omega\subset\mathbb{R}^{d}$ defined on $[0,T]\times\Omega\subset\mathbb{R}^{1+d}$ where $\Omega$ is a bounded domain in $\mathbb{R}^{d}$ with regular enough boundary $\partial\Omega$, the general form of a parabolic PDE is,
-$\mathbf{u}_{t}+{\mathcal{F}}[\mathbf{u}]=\mathbf{f}\,,$ (1)
-such that ${\mathcal{F}}[\cdot]$ is a linear or nonlinear differential operator, and $\mathbf{u}(t,\mathbf{x})$ denotes a unknown solution.
-The general initial and boundary conditions can be then formulated as:
-$\mathbf{u}(0,\mathbf{x})=\mathbf{g}(\mathbf{x}),\quad\mathbf{x}\in\Omega\,,$ (2)
-$\mathcal{B}[\mathbf{u}]=0,\quad t\in[0,T],\ \mathbf{x}\in\partial\Omega\,.$ (3)
-Here, $\mathbf{f}$ and $\mathbf{g}(\mathbf{x})$ are given functions with certain regularity; $\mathcal{B}[\cdot]$ denotes an abstract boundary operator, representing various boundary conditions such as Dirichlet, Neumann, Robin, and periodic conditions.
-We aim at approximating the unknown solution $\mathbf{u}(t,\mathbf{x})$ by a deep neural network $\mathbf{u}_{\mathbf{\theta}}(t,\mathbf{x})$, where $\mathbf{\theta}$ denotes the set of all trainable parameters of the network (e.g., weights and biases).
-If a smooth activation function is employed, $\mathbf{u}_{\mathbf{\theta}}$ provides a smooth representation that can be queried for any $(t,\mathbf{x})$.
+Physics-Informed Neural Networks.
+Given the spatio-temporal domain $\Omega\subset\mathbb{R}^{d}$ with spatial dimension $d$, defined on $[0,T]\times\Omega\subset\mathbb{R}^{1+d}$ where $\Omega$ is a bounded domain in $\mathbb{R}^{d}$ with boundary $\partial\Omega$, the general form of a parabolic PDE is, $u_{t}+{\mathcal{F}}[u]=f$, such that ${\mathcal{F}}[\cdot]$ is a linear or nonlinear differential operator, and $u(t,\mathbf{x})$ denotes a unknown solution.
+The general initial conditions are $u(0,\mathbf{x})=g(\mathbf{x})$, for $\mathbf{x}\in\Omega$, a boundary conditions $\mathcal{B}[\mathbf{u}]=0$, for $\quad t\in[0,T]$ and $\mathbf{x}\in\partial\Omega$.
+Here, $f$ and $g(\mathbf{x})$ are given and $\mathcal{B}[\cdot]$ denotes an abstract boundary operator.
+The general formulation of PINNs Raissi et al.
+(2019a) aim to approximate the unknown solution $u(t,\mathbf{x})$ by a deep neural network $u_{\theta}(t,\mathbf{x})$, where $\theta$ denotes the set of all trainable parameters of the network.
 The PDE residuals are defined as,
-$\mathcal{R}_{{\rm int}}[\mathbf{u}_{\mathbf{\theta}}](t,\mathbf{x})=\frac{\partial\mathbf{u}_{\mathbf{\theta}}}{\partial t}(t,\mathbf{x})+\mathcal{D}[\mathbf{u}_{\mathbf{\theta}}](t,\mathbf{x})-\mathbf{f}(\mathbf{x})\,,\quad(t,\mathbf{x})\in[0,T]\times\Omega\,,$ (4)
-and spatial and temporal boundary residuals, respectively, by
-$\mathcal{R}_{{\rm bc}}[\mathbf{u}_{\mathbf{\theta}}](t,\mathbf{x})=\mathcal{B}[\mathbf{u}_{\theta}](t,\mathbf{x})\,,\quad(t,\mathbf{x})\in[0,T]\times\partial\Omega\,,$ (5)
-and
-$\mathcal{R}_{{\rm ic}}[\mathbf{u}_{\mathbf{\theta}}](\mathbf{x})=\mathbf{u}_{\theta}(0,\mathbf{x})-\mathbf{g}(\mathbf{x})\,,\quad\mathbf{x}\in\Omega\,.$ (6)
+$\mathcal{R}_{r}[u_{\mathbf{\theta}}](t,\mathbf{x})=\frac{\partial u_{\mathbf{\theta}}}{\partial t}(t,\mathbf{x})+\mathcal{F}[u_{\mathbf{\theta}}](t,\mathbf{x})-f(\mathbf{x})\,.$ (1)
+The boundary and initial condition residuals are,
+$\mathcal{R}_{b}[u_{\theta}](t,\mathbf{x})$ $=\mathcal{B}[u_{\theta}](t,\mathbf{x})\,,$ (2)
+$\mathcal{R}_{i}[u_{\theta}](\mathbf{x})$ $=u_{\theta}(0,\mathbf{x})-g(\mathbf{x})\,.$ (3)
 Then, we train a physics-informed model by minimizing the following composite empirical loss:
-$\mathcal{L}(\mathbf{\theta}):=\underbrace{\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left|\mathcal{R}_{{\rm ic}}[\mathbf{u}_{\mathbf{\theta}}](\mathbf{x}_{ic}^{i})\right|^{2}}_{\mathcal{L}_{ic}(\mathbf{\theta})}+\underbrace{\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left|\mathcal{R}_{{\rm bc}}[\mathbf{u}_{\mathbf{\theta}}](t_{bc}^{i},\mathbf{x}_{bc}^{i})\right|^{2}}_{\mathcal{L}_{bc}(\mathbf{\theta})}+\underbrace{\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left|\mathcal{R}_{\rm int}[\mathbf{u}_{\mathbf{\theta}}](t^{i}_{r},\mathbf{x}^{i}_{r})\right|^{2}}_{\mathcal{L}_{r}(\mathbf{\theta})}\,,$ (7)
-which aims to enforce the neural network function $\mathbf{u}_{\theta}$ to satisfy the PDEs (1) with initial and spatial boundary conditions (2)–(3).
-The training data points $\{\mathbf{x}_{ic}^{i}\}_{i=1}^{N_{ic}}$, $\{t_{bc}^{i},\mathbf{x}_{bc}^{i}\}_{i=1}^{N_{bc}}$ and $\{t_{r}^{i},\mathbf{x}_{r}^{i}\}_{i=1}^{N_{r}}$ can be the vertices of a fixed mesh or points randomly sampled at each iteration of a gradient descent algorithm.
-2.2 Radial Basis Function Networks
-The general form of an RBF network for the training data points $\mathbf{x}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ with corresponding function values $y_{i}$ is given by,
-$f({\bf x})$ $=\sum_{j=1}^{N_{c}}c_{j}\phi(||{\bf x}-{\bf x}_{j}^{c}||),$ (8)
-where ${\bf x}^{c}=\{{\bf x}_{j}^{c}\}_{j=1}^{N_{c}}$ are the center points of the radial basis functions (RBFs).
-Here, $\phi$ represents the radial basis function, which are commonly chosen as the Gaussian, multi-quadric, or inverse multi-quadric functions, each offering different characteristics and benefits.
-The elements of the kernel matrix are given by:
-$\mathbf{K}_{ij}$ $=\phi(||{\bf x}_{i}-{\bf x}^{c}_{j}||),$ (9)
-where ${\bf x}_{i}$ represent the data points, and (9) evaluates the RBF based on the distance between the points.
-The kernel matrix captures the pairwise interactions between data points through the chosen RBF, and its structure is typically symmetric and positive-definite, assuming an appropriate choice of $\phi$.
-The choice of center points can vary depending on the application; in some cases, the center points may coincide with the data points, but they may also be selected independently of the data points to optimize approximation accuracy or computational efficiency.
-The RBFs determine how much influence each center has on the input, based on the distance between the input and the center.
+$\mathcal{L}(\theta):=\frac{\lambda_{ic}}{N_{ic}}\sum_{i=1}^{N_{ic}}\left|\mathcal{R}_{{\rm ic}}[u_{\theta}](\mathbf{x}_{ic}^{i})\right|^{2}$ (4)
+$+\frac{\lambda_{b}}{N_{b}}\sum_{i=1}^{N_{b}}\left|\mathcal{R}_{b}[u_{\theta}](t_{bc}^{i},\mathbf{x}_{b}^{i})\right|^{2}$ $+\frac{\lambda_{r}}{N_{r}}\sum_{i=1}^{N_{r}}\left|\mathcal{R}_{r}[u_{\theta}](t^{i}_{r},\mathbf{x}^{i}_{r})\right|^{2}\,,$ (5)
+which aims to enforce the neural network function $u_{\theta}$ to satisfy the PDE with initial and boundary conditions.
+$\lambda_{r}$, $\lambda_{b}$, and $\lambda_{i}$ are typically hyperparameters that control the interplay between the different loss terms.
+The training data points $\{\mathbf{x}_{ic}^{i}\}_{i=1}^{N_{ic}}$, $\{t_{bc}^{i},\mathbf{x}_{bc}^{i}\}_{i=1}^{N_{bc}}$ and $\{t_{r}^{i},\mathbf{x}_{r}^{i}\}_{i=1}^{N_{r}}$ can be fixed mesh or points randomly sampled at each iteration of a gradient descent algorithm.
+Radial Basis Function Networks.
+The general form of a radial basis function (RBF) network for a set of training data points $\mathbf{x}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ with corresponding function values $y_{i}$ is given by
+$f(\mathbf{x})$ $=\sum_{j=1}^{N_{c}}c_{j}\psi(||\mathbf{x}-\mathbf{x}_{j}^{c}||),$ (6)
+where $\mathbf{x}^{c}=\{\mathbf{x}_{j}^{c}\}_{j=1}^{N_{c}}$ represents the center points of the radial basis functions (RBFs), and $\psi$ is the chosen radial basis function.
+Common choices for $\psi$ include the Gaussian, multi-quadric, and inverse multi-quadric functions, each offering different approximation properties.
+The kernel matrix $\mathbf{K}\in\mathbb{R}^{N\times N_{c}}$ is constructed as $K_{ij}=\psi(||\mathbf{x}_{i}-\mathbf{x}_{j}^{c}||)$, where $\mathbf{x}_{i}$ represents the data points, and $K_{ij}$ evaluates the RBF response based on the distance between input points and center locations.
+The kernel matrix captures pairwise interactions between data points and is typically symmetric and positive-definite, assuming an appropriate choice of $\psi$.
+The selection of center points $\mathbf{x}^{c}$ plays a crucial role in the model’s performance.
+The RBFs determine the influence of each center on the input space based on the distance between $\mathbf{x}$ and $\mathbf{x}_{j}^{c}$.
+Gradient-based optimization is sometimes used to update both the center locations $\mathbf{x}^{c}$ and the RBF coefficients $c_{j}$ during training.
+This adaptive learning process defines an RBF network, where the radial basis functions act as the hidden layer activations, and the RBF coefficients serve as the trainable network parameters.
+The network is typically trained by minimizing the mean squared error (MSE) between the predicted function values and the given training data.
 […]
-Figure 1: Illustration of the RBF+PINN hybrid residual block with trainable strength connections between the RBF and PINN outputs.
-Gradient descent is sometimes used to optimize both the centers and the unknown weights ($c_{j}$) simultaneously.
-These approaches are referred to as RBF networks, where the radial basis functions are the hidden layer activations, and the RBF coefficients are the trainable network parameters.
-The mean-squared error between the linear combination of the hidden layer outputs and the known functions values is commonly minimized.
+Figure 1: Illustration of the hybrid residual block with trainable strength connections between the RBF and NN outputs.
 3 Hybrid Residual PINNs (HyResPINNs)
-In this section, we describe our proposed architecture—HyResPINNs. HyResPINNs  are a novel type of residual network such that each residual block incorporates an RBF kernel along with a standard DNN.
-3.1 Hybrid Residual Blocks
-Figure 1 shows the hybrid residual block specifics.
-We denote the output of the $l$-th residual block for the input ${\bf x}^{(l)}$ as $H^{(l)}({\bf x}^{(l)})$.
-Formally, the forward pass of each hybrid residual block in the HyResPINN  architecture is,
-$H^{(l)}({\bf x}^{(l)})=\sigma\big{(}\phi(\alpha^{(l)})F_{R}^{(l)}({\bf x}^{(l)})+(1-\phi(\alpha^{(l)}))F_{N}^{(l)}({\bf x}^{(l)})\big{)},$ (10)
-such that $\sigma$ is some non-linear activation function, $\alpha^{(l)}\in\mathbb{R}$ is a trainable parameter, $\phi$ represents the sigmoid function, $F_{R}^{(l)}({\bf x}^{(l)})$ is the output of the RBF network and $F_{N}^{(l)}({\bf x}^{(l)})$ is the output from a DNN or PINN within the $l$-th block.
-[…]
-Figure 2: Illustration of the HyResPINN  architecture using three blocks.
-Adaptive Hybrid Contribution Coefficients
-The trainable parameter $\alpha^{(l)}$ controls the relative contributions of the DNN and RBF components to the output—where optimizing $\alpha$ finds the best balance between these two elements.
-Since the sigmoid function $\phi$ constrains its output to the range $[0,1]$, the resulting output is a convex combination of the RBF and DNN components.
-We further incorporate adaptable residual connection parameters between each hybrid block, denoted as $\beta^{(l)}$—similar to the approaches in (Howard et al., 2023) and (Wang et al., 2024a)
-However, unlike (Wang et al., 2024a), where their adaptable parameters are initialized to zero to force the network to learn non-linearities from scratch, we initialize each $\alpha^{(l)}=0.5$, ensuring equal contributions from both components at the start of training, and each $\beta^{(l)}=1$.
-When $\alpha=1$, only the RBF network contributes; when $\alpha=0$, only the DNN contributes.
-The optimal choice of $\alpha$ depends on the problem characteristics.
-For problems with large regions of smoothness, the model might favor a lower $\alpha$, assigning more weight to the DNN components in each residual block.
-Conversely, for problems involving sharp transitions or discontinuities, a higher $\alpha$ (favoring the RBF) may be preferable.
-In hybrid problems with smooth and non-smooth regions, $\alpha$ will likely fall between 0.4 and 0.6, providing a balanced combination of the two networks.
-To encourage smoother solutions and prevent the network from introducing excessive non-linearity, we add a regularization term to each block’s trainable parameter $\alpha$.
-This regularization penalizes large values of $\alpha$, effectively controlling the contribution of non-linearity from the RBF components.
-Specifically, the total loss function is defined as:
-$\mathcal{L}=\lambda_{ic}\mathcal{L}_{ic}({\boldsymbol{\theta}})+\lambda_{bc}\mathcal{L}_{bc}({\boldsymbol{\theta}})+\lambda_{r}\mathcal{L}_{r}({\boldsymbol{\theta}})+\lambda_{p}\sum_{i=1}^{N_{blocks}}\alpha_{i}^{2},$ (11)
-where $\mathcal{L}_{ic}$ and $\mathcal{L}_{bc}$ represent the loss terms for the initial and boundary conditions, $\mathcal{L}_{r}$ represents the residual loss, and the final term $\lambda_{p}\sum_{i=1}^{N_{blocks}}\alpha_{i}^{2}$ acts as an $L_{2}$ regularization on the $\alpha$ parameters.
-The regularization helps balance the contributions of the smooth neural network and non-linear RBF components, promoting smoother solutions and more stable training.
-Figure 3: 2D Allen-Cahn equation: Comparison of (a subset of) the learned RBF kernels for a standard RBF network (top row) and the proposed hybrid RBF+NN residual block approach (bottom row).
-Each subplot shows the RBF value corresponding to different learned RBF centers in the input domain, marked by red crosses.
-Adaptive RBF Block Kernel
-The output of the RBF kernel $F_{R}^{(l)}({\bf x}^{(l)})$ can be described as,
-$F_{R}^{(l)}({\bf x}^{(l)})=K({\bf x})W,$ (12)
-such that the kernel matrix $K$ is defined as (9) where the kernel matrix $K$ is based on the Wendland $C^{4}$ kernel in 2D. In this work, we focus on isotropic version of the Wendland kernel.
-The isotropic Wendland $C^{4}$ kernel for input ${\bf x}$ and the $i$-th center is defined as:
-$\phi_{i}(\mathbf{x})=\left(1-\frac{\|{\bf x}-{\bf x}^{c}_{i}\|}{\tau_{i}}\right)_{+}^{6}\left(35\left(\frac{\|{\bf x}-{\bf x}^{c}_{i}\|}{\tau_{i}}\right)^{2}+18\frac{\|{\bf x}-{\bf x}^{c}_{i}\|}{\tau_{i}}+3\right),$ (13)
-where ${\bf x}^{c}_{i}$ is the center of the $i$-th RBF, $\tau$ is a scaling parameter, and $\|{\bf x}-{\bf x}^{c}_{i}\|$ represents the Euclidean distance between ${\bf x}$ and ${\bf x}^{c}_{i}$.
-The Wendland kernel is compactly supported, meaning $\phi_{i}(\mathbf{x})=0$ for $\|{\bf x}-{\bf x}^{c}_{i}\|\geq\tau_{i}$, which leads to sparse kernel matrices and computational efficiency.
-Each $\tau_{i}$ is a trainable parameter in the network and optimize through gradient descent along with all other network parameters.
-Figure 3 demonstrates learned scale parameters within an optimized RBF kernel.
-Block Neural Network
-Following the convention of Cyr et al.
-(2020), we represent the output of the DNN $F_{N}^{(l)}({\bf x}^{(l)})$ $F_{N}^{(l)}\in\mathbb{R}^{d}\rightarrow\mathbb{R}$ of width $w$, as a linear combination of adaptive basis functions given by
-$F_{N}^{(l)}({\bf x};a,\theta^{h})=\sum_{i=1}^{w}a_{i}\sigma_{i}{\left({\bf x};\theta^{h}\right)},$ (14)
-where each $a_{j}$ for $j=1,..,w$ and $\theta^{h}$ constitute the weights and biases in the last layer and hidden layers respectively, forming the set of all network parameters $\theta$.
-Then, each $\sigma_{j}$ are non-linear smooth activation functions such as Tanh acting on the outputs of the hidden layers.
-We choose to use a standard DNN architecture, but any DNN architecture such as ResNets, would work.
-The parameters $\theta$ are computed through some iterative optimization technique.
-In this work, we use variants of gradient descent methods such as ADAM Kingma & Ba (2015) and L-BFGS Liu & Nocedal (1989).
-Figure 2 shows a visualization of our full model architecture with adaptive residual block skip connections, along with the input block structure used to lift the inputs to the desired higher dimension, and the output neural network block used to project each block’s output down to the output dimension.
+We propose HyResPINN, a deep learning framework that integrates radial basis function networks (RBFs) and feedforward neural networks (NNs) within a hybrid residual structure.
+This architecture enhances function approximation by leveraging localized adaptivity from RBFs and global smoothness from NNs.
+Hybrid Block $1$Hybrid Block $2$ Hybrid Block $3$ $\mathbf{x}$$\bar{\mathbf{a}}^{(1)}$$\bar{\mathbf{a}}^{(2)}$$\bar{\mathbf{a}}^{(3)}$$\mathbf{y}$$\phi(\beta^{(1)})$$\phi(\beta^{(2)})$$\phi(\beta^{(3)})$
+Figure 2: Illustration of the HyResPINN architecture using three hybrid residual blocks. The trainable skip connections $\phi(\beta^{(l)})$ modulate the information flow between blocks.
+3.1 Hybrid Residual Block Structure
+The core component of HyResPINN is the hybrid residual block, illustrated in Figure 1.
+Each block consists of two primary components: an RBF-Net, denoted as $F_{R}$, which captures localized behavior via an RBF kernel matrix, and an NN component, denoted as $F_{N}$, which provides a globally smooth representation of the solution.
+These two outputs are combined using a trainable weighting function $\phi(\alpha^{(l)})$, which dynamically adjusts the contributions of the RBF-Net and NN within each block.
+Given a set of $N$ points, $\mathbf{x}\in\mathbb{R}^{N\times d}$, the output of the $l$-th hybrid block is computed as:
+$F^{(l)}(\mathbf{x})=\phi(\alpha^{(l)})F_{R}^{(l)}(\mathbf{x})+\left(1-\phi(\alpha^{(l)})\right)F_{N}^{(l)}(\mathbf{x}),$ (7)
+where $F_{R}^{(l)}(\mathbf{x})\in\mathbb{R}^{N\times p}$ represents the output of the RBF-Net, while $F_{N}^{(l)}(\mathbf{x})\in\mathbb{R}^{N\times p}$ corresponds to the NN output.
+Similar to Howard et al.
+(2023), the function $\phi(\alpha^{(l)})$ is a learnable gating mechanism that determines the relative contributions of the two models.
+To introduce non-linearity between block outputs and constrain them within a well-defined range, each block output $F^{(l)}$ is passed through a nonlinear activation function $\sigma(\cdot)$.
+In this work, $\sigma=\tanh$ is chosen to ensure that the outputs remain within the bounded space $[-1,1]^{d}$.
+RBF Block Component.
+The RBF-Net component is defined by the function:
+$F_{R}^{(l)}(\mathbf{x})=W^{(l)}\psi(||\mathbf{x}-\mathbf{x}_{j}^{c}||),$ (8)
+where $W^{(l)}\in\mathbb{R}^{N\times p}$ are the trainable RBF coefficients, $\mathbf{x}_{j}^{c}$ represents the RBF centers, and $\psi(||\mathbf{x}-\mathbf{x}_{j}^{c}||)$ is a radial basis function that applies a distance-based transformation to the input $\mathbf{x}$.
+The output dimension of the block is denoted by $p$, which serves as a user-defined hyperparameter.
+This component provides localized adaptivity, allowing the network to capture fine details and sharp transitions in the function approximation.
+The choice of $\psi$ influences the behavior of the model, with common options including Gaussian, multi-quadric, and inverse multi-quadric functions.
+Wendland Kernel Selection.
+In this work, we focus on an isotropic Wendland $C^{4}$ kernel, which is compactly supported, ensuring both sparsity and computational efficiency. The Wendland kernel is defined as:
+$\psi_{i}(\mathbf{x})$ $=\left(1-\frac{r_{i}}{\tau_{i}}\right)_{+}^{6}\left(35\left(\frac{r_{i}}{\tau_{i}}\right)^{2}+18\frac{r_{i}}{\tau_{i}}+3\right),$ (9)
+where $\mathbf{x}_{i}^{c}$ represents the center of the $i$-th RBF, $\tau_{i}$ is a trainable scaling parameter, and $r_{i}=\|\mathbf{x}-\mathbf{x}_{i}^{c}\|$ corresponds to the Euclidean distance between $\mathbf{x}$ and $\mathbf{x}_{i}^{c}$.
+Each $\tau_{i}$ is optimized through gradient descent, allowing the network to dynamically adapt the kernel shape throughout training.
+Figure 3 demonstrates a subset of the learned scale parameters for an optimized RBF kernel.
+Figure 3: 2D Allen-Cahn equation: Comparison of two random RBF kernels from a standard RBF-Net (top row) and the first block of a HyResPINNs (bottom row).
+Each subplot shows the (final) learned RBF scale parameter ($\tau$) for the selected center points (marked by red crosses).
+NN Block Component.
+The NN component is defined as:
+$F_{N}^{(l)}(\mathbf{x})=\mathcal{NN}(\mathbf{x};\theta),$ (10)
+where $\theta$ represents the set of trainable parameters in the network.
+The NN consists of at least one hidden layer with a non-linear activation function, which enables the model to approximate complex functions while maintaining smoothness.
+Unlike the RBF component, which specializes in localized features, the NN component is designed to capture global trends in the solution.
+Although a standard deep neural network is used in this work, other architectures such as residual networks (ResNets) could be employed in place of the NN component within the hybrid residual block.
+Adaptive Block Parameter.
+To balance the contributions of $F_{R}^{(l)}$ and $F_{N}^{(l)}$, we employ a trainable weighting function $\phi(\alpha^{(l)})$, where $\phi$ is the sigmoid function, ensuring that $\phi(\alpha^{(l)})$ remains within the range $[0,1]$.
+We initialize $\phi(\alpha^{(l)})=0.5$, ensuring that both components contribute equally at the start of training.
+The learnable parameter $\alpha^{(l)}$ dynamically adjusts throughout training to optimize the relative influence of the NN and RBF components based on the characteristics of the problem.
+3.2 Deep Hybrid Residual Architecture
+The full HyResPINN architecture, shown in Figure 2, consists of multiple hybrid residual blocks stacked within a residual learning framework.
+Each block’s output is propagated forward, with trainable skip connections modulated by $\phi(\beta^{(l)})$.
+The model computes the intermediate representations $\bar{\mathbf{a}}^{(l)}$ at each layer according to:
+$\bar{\mathbf{a}}^{(l)}=\phi(\beta^{(l)})F^{(l)}(\mathbf{x})+(1-\phi(\beta^{(l)}))\bar{\mathbf{a}}^{(l-1)}.$ (11)
+After passing through $L$ hybrid residual blocks, the final prediction is obtained via a deep neural network (DNN), which refines the hybrid representation into the final output:
+$\mathbf{y}=\text{DNN}(\bar{\mathbf{a}}^{(L)}).$ (12)
+Unlike PirateNets (Wang et al., 2024), which initialize their residual connection parameter to zero at the start of training, HyResPINN initializes each $\phi(\beta^{(l)})=1$.
+This initialization ensures that the early training dynamics do not force the network to learn non-linearities from scratch, but rather begin with a meaningful blend of residual and block outputs.
+PirateNets also use an additive residual formulation rather than the convex combination approach in HyResPINN, which allows our method to explicitly control the contribution of each block’s output through a learned weighting function.
+This structured balancing mechanism provides greater flexibility in adapting to both smooth and non-smooth solution structures.
+Regularization.
+To encourage smoothness and prevent excessive non-linearity, we introduce an $L_{2}$ penalty on the trainable parameters $\alpha$ and $\beta$:
+$\mathcal{L}$ $=\lambda_{i}\mathcal{L}_{i}+\lambda_{b}\mathcal{L}_{b}+\lambda_{r}\mathcal{L}_{r}+\lambda_{p}\sum_{i=1}^{N_{blocks}}\left(\alpha_{i}^{2}+\beta_{i}^{2}\right).$ (13)
+This regularization term helps stabilize the training process by discouraging excessive reliance on either the NN or RBF component in each block.
+By contrast, PirateNets do not apply explicit regularization to their trainable residual parameters, which can lead to less controlled adaptation between blocks.
+The additional regularization in HyResPINN ensures that the learned weighting functions $\phi(\alpha^{(l)})$ and $\phi(\beta^{(l)})$ remain well-behaved throughout training, preventing overfitting to sharp local variations and promoting a more balanced function approximation.
 Problem Domain Boundary Cond. PINN ResPINN Expert Stacked PirateNet HyResPINN
@@
 (rough coefficients) 2D Box Neumann $6.85\times 10^{-5}$ $2.69\times 10^{-5}$ $0.00011$ $0.00015$ $5.44\times 10^{-5}$ $\mathbf{1.05\times 10^{-5}}$
-Table 1: Relative $L^{2}$ test error results for various PDE problems and baseline methods.
+Table 1: Overview of the relative $L^{2}$ test error results for various PDE problems and baseline methods.
 4 Experimental Results and Discussion
 In this section, we demonstrate the effectiveness of the proposed HyResPINNs  architecture across a diverse collection of benchmark problems compared to the leading baseline methods.
-We show that HyResPINNs  consistently produce lower relative $L^{2}$ errors between the predicted and ground truth solutions for the same training set size when compared to the applicable baselines trained under the same experimental procedures and hyperparameters.
-We compare each method on the 1D non-linear hyperbolic Allen-Cahn equation and the Darcy Flow equation in two and three dimensions with both Dirichlet and Neumann boundary conditions.
-Our main results are summarized in Table 1.
-Figure 4: 2D Allen-Cahn equation: Comparison of the predicted solutions for the Allen-Cahn equation using HyResPINN, standard PINN, and RBF network models.
-The top row shows the predicted solutions at for HyResPINN (left), standard PINN (center), and RBF network (right).
-The second row illustrates the absolute error between the predicted solutions and the exact solution.
-The bottom row shows the predicted solutions for three time steps comparing the exact solution (green), HyResPINN (orange), and standard PINN (dashed blue).
-The HyResPINN model effectively captures both the smooth and sharp features of the solution, while the standard PINN struggles with sharp transitions, leading to larger errors in these regions.
+We first tested each method on the 1D non-linear hyperbolic Allen-Cahn equation to examine how the local RBFs in the HyResPINN  blocks help to predict sharp solution transitions.
+We then tested the Darcy Flow equation in 2D and 3D annulus/extruded annulus domains (respectively) with Neumann and Dirichlet boundary conditions to assess how HyResPINNs predict solutions on irregular domains, evaluating its robustness in handling geometric complexities.
+Next, we examined the 2D Darcy Flow problem with rough coefficients to evaluate how our method handles discontinuities and non-smooth solutions.
+Finally, we tested the Kuramoto–Sivashinsky equation to test our method on regular and chaotic problems, demonstrating its capability to model complex spatiotemporal dynamics.
+Appendix B details all PDE benchmarks and corresponding experimental settings.
+Table 1 summarizes our main results.
+Figure 4: Allen-Cahn equation: Comparison of the predicted solutions for the Allen-Cahn equation using HyResPINN, standard PINN, and RBF network models.
+The top row shows the predicted solutions for HyResPINN (left), standard PINN (center), and RBF PINN (right).
+The second row shows the absolute error between the predicted and true solutions.
+The bottom row shows the predicted solutions for time steps ($t=0.25,0.5,0.99$) compared to the true solution.
 Baseline Methods.
-We compare the proposed HyResPINN  against a suite of baseline methods under exactly the same hyper-parameter settings in each example.
-Baselines include the standard PINN as originally formulated in (Raissi et al., 2019) (PINN), a PINN based on the architectural guidelines proposed in Wang et al.
-(2023) (ExpertPINNs), PINNs with residual connections (ResPINNs), PirateNets Wang et al.
-(2024a), and the stacked PINN approach of (Howard et al., 2023) (StackedPINNs).
-To ensure a fair comparison, we implemented the approaches detailed in (Raissi et al., 2019; Wang et al., 2024a; 2023; Howard et al., 2023) following the architectural details provided in each.
+We compare the proposed HyResPINN  against a suite of baseline methods, including 1) the standard PINN as originally formulated in (Raissi et al., 2019b) (PINN), 2) a PINN based on the architectural guidelines proposed in Wang et al.
+(2023) (ExpertPINNs), 3) PINNs with residual connections (ResPINNs), 4) PirateNets Wang et al.
+(2024), and 5) the stacked PINN approach of (Howard et al., 2023) (StackedPINNs).
+To ensure a fair comparison, we implemented the approaches detailed in (Raissi et al., 2019b; Wang et al., 2024; 2023; Howard et al., 2023) following the architectural details provided in each.
 Experimental Setup.
 We follow similar experimental design procedures as those described in Wang et al.
-(2022a; 2023; 2024a).
-A full description of the experimental setups are included in Appendix A, and exact hyper-parameter setting are detailed in Table 2.
-We train all models on Nvidia A100 GPU running Centos 7.2.
-Code for our methods and all compared baseline approaches is written in libtorch $1.11$ (the C++ version of PyTorch) and will be made publicly available upon publication.
-4.1 1D Allen-Cahn
-We first focus on the Allen-Cahn equation, a challenging benchmark for conventional PINN models that has been extensively studied in recent literature  (Wight & Zhao, 2020; Wang et al., 2022a; Daw et al., ).
+(2022a; 2023; 2024).
+The code for our methods and all compared baseline approaches is written in LibTorch $1.11$ (the C++ version of PyTorch) and will be publicly available upon publication.
+We train all models on a Nvidia A100 GPU running Centos 7.2.
+A full description of the experimental setups are included in Appendix C.
+Figure 5: Allen-Cahn equation: Comparison of the mean relative $L^{2}$ error using various methods as a function of the number of hidden layers.
+4.1 Allen-Cahn
+First, we present results for 1D the Allen-Cahn problem with periodic boundary conditions, a well-known challenge for conventional PINN models and a widely studied benchmark in recent literature  (Wight & Zhao, 2020; Wang et al., 2022a; Daw et al., ).
+This problem is particularly difficult for two reasons.
+First, the Allen-Cahn PDE requires PINNs to approximate solutions with sharp transitions in space and time, posing challenges for standard neural network architectures.
+Second, there is an inherent incompatibility between the initial and boundary conditions, leading to a weak discontinuity at time $t=0$.
+Traditional numerical methods would likely produce oscillations near the discontinuity due to the mismatch between the initial and boundary conditions.
+However, unlike traditional solvers, PINNs do not rely on discretization and instead learn their model parameters through optimization.
+This optimization process may mitigate spurious oscillations, as it relies on stochastic updates and continuous function approximations.
+Specifically, implicit regularization from both mini-batch training and stochastic gradient descent (SGD) may contribute to stabilizing optimization and reducing any spurious oscillations Sekhari et al.
+(2021).
+Therefore, despite this discontinuity, we evaluate our method on this benchmark due to its established role in assessing PINN performance.
+Hybrid Residual Blocks Improve Prediction Accuracy.
+Figure 4 demonstrates the advantages of the hybrid residual block structure in capturing both smooth and sharp features in the 1D Allen-Cahn solution.
+The absolute error plots show that HyResPINN exhibits significantly lower errors than standard PINNs and RBF-PINNs, validating its improved accuracy.
+As expected, the PINN prediction provides a smooth approximation but struggles with sharp transitions due to the inherent smoothness of neural networks.
+Meanwhile, RBF-Net fails to recover both the global structure and fine local details because, while its length scale is learned through optimization, its center locations and number of basis functions remain fixed.
+This rigid structure prevents it from dynamically placing basis functions in regions of rapid change, leading to poor approximation of sharp features.
+HyResPINN resolves these limitations by combining a smooth neural network with localized RBF adaptivity.
+The neural network component ensures global consistency, while the RBF kernels allow for finer, more flexible local approximations.
+This synergy enables the model to balance smooth and non-smooth behavior, significantly improving predictive accuracy.
+Figure 6: Allen-Cahn equation: Comparison of mean relative $L^{2}$ error across various methods, plotted against the number of training iterations (left) and the mean training time (right).
+Vertical bars in right plot indicate the time for which each method achieved an error of $10^{-2}$.
+Depth in Residual Blocks Improves Accuracy.
+Figure 5 demonstrates that increasing the number of residual blocks in HyResPINN leads to consistently lower errors, outperforming the competing methods.
+Residual learning allows deeper networks to refine predictions iteratively, progressively improving both global structure and sharp transitions.
+The competing methods struggle to maintain accuracy as depth increases, whereas HyResPINN remains robust, demonstrating the effectiveness of deep hybrid residual learning.
+Figure 7: 2D and 3D Smooth Darcy Flow equation: Comparison of the mean relative $L^{2}$ errors across each baseline method for the 2D Darcy Flow problem (left) and the 3D Darcy flow problem (right) plotted against the number of training collocation points.
+Solid lines show the solution error, while dashed lines show the x-directional flux errors.
+Training Efficiency and Convergence of HyResPINNs.
+Figure 6 (left) shows that HyResPINN achieve the lowest error over iterations, demonstrating the most consistent reduction in error.
+This suggests that the hybrid residual structure efficiently utilizes depth to refine solutions progressively.
+While the right plot in Figure 6 shows that HyResPINN has a slightly higher training cost than other methods, its significantly lower error justifies this cost.
+In particular, for long training times, HyResPINN consistently outperforms all baselines, highlighting its robustness in solving complex PDEs.
+The vertical bars indicate the training times required for each method to reach an error threshold of $10^{-2}$, further demonstrating HyResPINN’s efficiency in achieving high-accuracy solutions.
+This makes the training cost vs.
+accuracy trade-off more explicit and clearly separates training efficiency from accuracy discussions.
+4.2 Darcy Flow with Smooth Coefficients
+We next test the performance of HyResPINN on the Darcy Flow equation with smooth coefficients.
+Darcy Flow serves as a fundamental model for various physical processes, including porous media flow, heat transfer, and semiconductor doping.
+Since real-world applications often involve solving these equations under varying boundary conditions, it is essential to evaluate both solution accuracy and robustness to different constraints.
+To assess convergence, we construct smooth manufactured solutions in two settings: a 2D annular domain and a 3D cylindrical domain obtained by extruding the 2D geometry along the z-axis.
+The exact solutions are given as,
+$u(x,y)$ $=\sin(x)\sin(y),$ (14)
+$u(x,y,z)$ $=\sin(x)\sin(y)\sin(z),$ (15)
+for the 2D and 3D cases, respectively.
+By substituting these exact solutions into the model problem we define the source term and the boundary data.
+The true solutions are shown in Figure 9 for both the 2D and 3D cases.
+The detailed problem setup, including boundary conditions and parameter choices, is outlined in Appendix B.2.
+We perform a convergence study by analyzing the mean relative $L^{2}$ error as the number of training collocation points increases.
+The problem is solved under both Dirichlet and Neumann boundary conditions to test robustness.
+Figure 7 presents the results for both 2D and 3D cases, where solid lines represent the solution error, and dashed lines correspond to the x-directional flux errors.
+This experiment allows us to evaluate not only the accuracy of HyResPINN but also its ability to efficiently utilize training data.
+A robust method should achieve lower errors with fewer training points while remaining stable across different boundary conditions.
+The results in Figure 7 demonstrate that HyResPINN consistently achieves the lowest errors compared to baseline methods across all training sizes.
+Notably: HyResPINN achieves high accuracy with fewer training points, demonstrating its ability to generalize well with limited data.
+Unlike traditional PINN-based approaches that may exhibit sensitivity to Dirichlet or Neumann constraints, HyResPINN remains stable across different boundary types.
+The method extends seamlessly from 2D to 3D without significant degradation in performance, confirming its suitability for large-scale PDE problems.
+These results further validate the effectiveness of hybrid residual learning in solving elliptic PDEs with smooth solutions.
+4.3 Darcy Flow with Rough Coefficients
+Figure 8: 2D Rough Darcy Flow equation: Comparison of the mean relative $L^{2}$ errors across various methods, plotted against the number of training collocation points.
+Solid lines show the solution error, while dashed lines show the x-directional flux errors.
+A key challenge when solving the Darcy Flow equation 16 in scenarios with discontinuous coefficients is accurately approximating the flux variable, particularly its normal and tangential components.
+The normal component of the flux remains continuous across material interfaces, whereas the tangential component may exhibit discontinuities.
+This behavior is crucial for physically accurate solutions, especially when solving problems involving heterogeneous materials or rough coefficients.
+In contrast, traditional collocated methods—such as Galerkin, stabilized Galerkin, and least-squares finite elements—often struggle to capture these discontinuities in flux, typically resulting in oscillations at material interfaces.
+These oscillations indicate that the methods fail to accurately respect the physical continuity of the flux field.
+In this section, we demonstrate that HyResPINN and its variants effectively capture the physical behavior of flux, providing accurate approximations of both the normal and tangential components in the presence of discontinuous coefficients.
+This is particularly evident in the convergence analysis for the "five strip problem", a well-documented manufactured solution test Nakshatrala et al.
+(2006); Masud & Hughes (2002), which evaluates the ability of a method to preserve the continuity of normal flux across material interfaces.
+As shown in Figure 10, the true solutions for both the solution field and the flux components are depicted for reference.
+We solve the problem with Neumann boundary conditions, which are commonly used for this problem setup.
+Convergence results are presented in Figure 8, where we compare HyResPINN against baseline methods.
+The results show that HyResPINN is capable of accurately capturing the solution and flux, though it requires larger training sizes compared to the other methods.
+Nonetheless, it outperforms the baselines in terms of accuracy and physical fidelity, particularly in approximating the flux field without introducing undesirable oscillations.
+4.4 Kuramoto–Sivashinsky Equation
+In this section, we use the Kuramoto–Sivashinsky equation as a benchmark problem under periodic boundary conditions.
+The detailed problem setup is provided in Appendix B.
+The Kuramoto–Sivashinsky equation is a fourth-order nonlinear partial differential equation widely used to model spatiotemporal patterns in fluid dynamics, combustion, and other systems exhibiting chaotic behavior.
+Solving this equation allows us to assess HyResPINN’s ability to approximate highly complex and chaotic systems.
+HyResPINN achieved relative $L_{2}$ errors of $15.610\%$ for the regular case and $64.50\%$ for the chaotic case, while ExpertPINNs achieved significantly lower errors of $2.120\%$ and $3.272\%$, respectively.
+These results highlight the need for further refinement of HyResPINN, particularly in handling chaotic regimes.
+5 Conclusion
+In this work, we introduced HyResPINNs, a novel class of physics-informed neural networks that incorporate adaptive hybrid residual blocks combining the strengths of standard neural networks and radial basis function (RBF) networks.
+Our architecture effectively leverages smooth and non-smooth function approximators.
+The adaptive combination parameters within each block allow the model to balance the contributions of neural and RBF components during training.
+Our experiments demonstrate that HyResPINNs  outperform traditional PINNs and state-of-the-art methods in accuracy and robustness, particularly for problems involving mixed smooth and non-smooth regions.
+This work increases the flexibility and generalizability of PINNs by bridging the gap between classical numerical methods and DNN-based approaches for solving PDEs.
+Appendix A Computational Cost
+Problem Domain Boundary Cond. PINN ResPINN Expert Stacked PirateNet HyResPINN
+Allen-Cahn 1D Space/Time Periodic $34.99$ $17.65$ $64.81$ $22.29$ $116.72$ $150.88$
+DarcyFlow 2D Annulus Neumann $2.37$ $2.76$ $18.67$ $2.18$ $11.88$ $5.73$
+(smooth coefficients) Dirichlet $2.08$ $2.09$ $17.64$ $1.64$ $10.87$ $11.87$
+3D Annulus Neumann $8.1$ $8.5$ $27.6$ $6.9$ $20.8$ $35.9$
+Dirichlet $5.9$ $6.0$ $21.3$ $5.2$ $15.3$ $30.1$
+(rough coefficients) 2D Box Neumann $2.4$ $2.4$ $7.7$ $2.1$ $7.7$ $8.4$
+Table 2: Average wall-clock training time (in minutes) for each PDE problem and baseline methods.
+Appendix B Benchmark PDEs
+B.1 Allen-Cahn Equation
 For simplicity, we consider the one-dimensional case with a periodic boundary condition with $t\in[0,1]$, and $x\in[-1,1]$:
@@
 $u(t,-1)=u(t,1)\,,\;u_{x}(t,-1)=u_{x}(t,1)\,.$
-The Allen-Cahn PDE is an interesting benchmark for PINNs as it introduces periodic boundary conditions, and because it is a “stiff” PDE that challenges PINNs to approximate solutions with sharp space and time transitions.
-Hybrid residual blocks improve prediction accuracy.
-We first demonstrate that the hybrid residual block structure introduced in our method enhances the ability to capture both smooth and sharp features in the 2D Allen-Cahn solution.
-As shown in Figure 4, the proposed approach outperforms competing baselines, which struggle to accurately represent the sharp transitions in the solution.
-This difficulty arises from the smoothness constraints of standard neural network architectures.
-In contrast, our method effectively captures these sharp transitions by leveraging a combination of smooth neural networks and adaptive non-smooth RBF kernels, allowing the model to balance between global smooth behavior and localized discontinuities.
-Figure 3 shows a random subset of the learned RBF kernels captured by the proposed hybrid approach.
-Specifically, it captures similar RBF structures (to a standard RBF network), but benefits from the additional flexibility provided by the NN component, allowing for more nuanced function approximation in regions requiring a balance of smoothness and sharp transitions.
-This is likely apparent in the on average smaller smaller kernels learned by the hybrid block.
-Stacking structure of residual blocks improve prediction accuracy.
-Figure 5: Allen-Cahn equation: Comparison of the mean relative $L^{2}$ error using various methods as a function of the number of hidden layers.
-We next show that the stacking structure employed by the HyResPINN  in each hybrid residual block structure introduced here further helps to capture the smooth and sharp features in the Allen-Cahn solution.
-Figure 5 shows the predicted solutions between the proposed method and the competing baselines.
-The competing methods need help accurately capture the sharp transitions in the solutions.
-In contrast, our approach more accurately captures these sharp transitions due to the conjunction of the smooth neural network architecture and the learned non-smooth RBF kernels.
-Figure 6: Allen-Cahn equation: Comparison of mean relative $L^{2}$ error across various methods, plotted against the number of training iterations (left) and the mean training time (right).
-Vertical bars in right plot indicate the time for which each method achieved an error of $210^{-}$ 2.
-We next evaluate the effect of network depth on predictive accuracy for different architectures.
-Figure 5 HyResPINN consistently achieves the lowest error across different network depths.
-The performance of other methods shows varying sensitivity to the number of hidden layers, with HyResPINN offering the most robust performance.
-Further, the left plot in Figure 6 shows that HyResPINN and PiratePINN achieve the lowest error over iterations, with HyResPINN consistently reducing the error more effectively than the other methods.
-The right plot demonstrates that—although HyResPINN has a slightly higher training cost—it significantly outperforms other methods in accuracy, particularly for long training times.
-These results highlight the robustness and efficiency of HyResPINN in solving complex PDEs.
 Remark
-While attempting to replicate the results of (Wang et al., 2024a), using the reported hyperparameters and implementation details, we observed a slight performance gap between the results obtained in our implementation and those listed in their paper.
+While attempting to replicate the results of (Wang et al., 2024), using the reported hyperparameters and implementation details, we observed a slight performance gap between the results obtained in our implementation and those listed in their paper.
 The observed discrepancy in the final model’s performance could stem from several factors such as subtle variations in the implementation environment—such as hardware differences, software library versions, or random seed initialization.
-Despite the differences between our replication results and those reported in (Wang et al., 2024a), the comparison between our architecture and the replicated version of their approach remains fair, as both were tested under identical conditions using the same dataset, hyperparameters, and training setup (including hardware and software environment).
+Despite the differences between our replication results and those reported in (Wang et al., 2024), the comparison between our architecture and the replicated version of their approach remains fair, as both were tested under identical conditions using the same dataset, hyperparameters, and training setup (including hardware and software environment).
 Our findings indicate that our architecture outperforms the replicated version of their model in many cases, along with all other replicated baseline methods.
-4.2 Darcy Flow
+B.2 Darcy-Flow Equation
+In the Allen-Cahn experiment, we employed mini-batching with randomly sampled collocation points.
+However, for the Darcy Flow experiments, we aimed to systematically study the effect of increasing the number of collocation points on convergence.
+To achieve this, we used static training datasets, ensuring that all experiments maintained a fixed set of training points per run.
+This allows for a controlled evaluation of error reduction as the dataset size increased, as shown in Figure 7.
+The use of static datasets also eliminated variability from stochastic sampling, enabling clearer comparisons between problem dimensions and boundary conditions.
+This approach provides a more precise understanding of the trade-off between accuracy and the number of training points, highlighting the efficiency of HyResPINN in solving elliptic PDEs.
+B.2.1 Smooth Coefficients
+Figure 9: Smooth Darcy Flow equation: True solutions of the 2D and 3D problems.
 In this section, we describe the Darcy Flow problem, an elliptic boundary value problem given by,
@@
 While Darcy’s law governs fluid flow through porous media based on pressure gradients, the diffusion of dopants in semiconductors is guided by concentration gradients, analogous to how fluids move through porous materials in Darcy’s law.
-Figure 7: 2D and 3D Smooth Darcy Flow equation: Comparison of the mean relative $L^{2}$ errors across each baseline method for the 2D Darcy Flow problem (left) and the 3D Darcy flow problem (right) plotted against the number of training collocation points.
-Solid lines show the solution error, while dashed lines show the x-directional flux errors.
-4.2.1 Annulus Domain with Smooth Coefficients
-We initially showcase the convergence rate of our method by employing smooth manufactured solutions.
-We consider two scenarios: one within an annular 2D domain, and another within a 3D domain formed by extruding the 2D geometry in the z-direction, resulting in a cylinder with a height of two.
-We present a convergence study analyzing the error rate as the number of points increases.
-The exact solutions are set as,
-$u(x,y)=\sin(x)\sin(y),\;\;\;\;u(x,y,z)=\sin(x)\sin(y)\sin(z),$ (16)
-for the 2D and 3D cases, respectively.
-By substituting these exact solutions into the model problem we define the source term and the boundary data.
-We solve the discrete problem with either Dirichlet or Neumann boundary conditions.
-Convergence results are shown in Figure 7 for the two and three dimensional problems and demonstrate that HyResPINNs  remain robust to boundary condition type, size of training set (achieves lowest errors compared to the baselines for fewer training points), and problem dimension.
-4.2.2 2D Box Domain with Rough Coefficients
-A defining feature of methods for solving the Darcy Flow equation 15, lies in their proficiency in accurately depicting the flux variable in scenarios where the coefficients $\mu$ are discontinuous.
-In these instances, the flux’s normal component maintains continuity across material interfaces, whereas the tangential component may exhibit discontinuities.
-In contrast, collocated methods like Galerkin, stabilized Galerkin, and least-squares finite elements often fail to replicate this physical behavior, typically resulting in oscillations at the interface.
-In this section, we demonstrate that HyResPINNs  successfully provides physically accurate flux approximations for problems characterized by discontinuous coefficients.
-The initial example presented is the well-documented "five strip problem" Nakshatrala et al.
-(2006); Masud & Hughes (2002), which serves as a conventional manufactured solution test to evaluate a method’s capability to preserve the continuity of normal flux.
+B.2.2 Rough Coefficients
+Figure 10: (Left) True solution on the 2D box domain. (Middle) True transverse flux. (Right) True longitudinal flux.
 The prescribed exact solution on domain $\Omega=[0,1]^{2}$ is given by,
@@
 with different $\mu_{i}$ on each $\Omega_{i}$ such that $\mu_{1}=16,\;\;\mu_{2}=6,\;\;\mu_{3}=1,\;\;\mu_{4}=10,\;\;\mu_{5}=2$.
-We solve the discrete problem with Neumann boundary conditions.
-Convergence results are shown in Figure 8 and show that HyResPINNs  provide accurate solutions, albeit with larger training sizes.
-Figure 8: 2D Rough Darcy Flow equation: Comparison of the mean relative $L^{2}$ errors across various methods, plotted against the number of training collocation points.
-Solid lines show the solution error, while dashed lines show the x-directional flux errors.
-5 Conclusion
-In this work, we introduced HyResPINNs, a novel class of physics-informed neural networks that incorporate adaptive hybrid residual blocks combining the strengths of standard neural networks and radial basis function (RBF) networks.
-Our architecture effectively captures continuous and discontinuous features by leveraging smooth and non-smooth function approximators.
-The adaptive combination parameters within each block allow the model to balance the contributions of neural and RBF components during training.
-Furthermore, including Wendland kernels enhances the model’s ability to handle sharp transitions while maintaining computational efficiency.
-Our experiments demonstrate that HyResPINNs  outperform traditional PINNs and state-of-the-art methods in accuracy and robustness, particularly for problems involving mixed smooth and non-smooth regions.
-This work increases the flexibility and generalizability of PINNs by bridging the gap between classical numerical methods and DNN-based approaches for solving PDEs.
-Appendix A Experimental Design
+B.3 Kuramoto–Sivashinsky Equation
+We use the 1D Kuramoto–Sivashinsky equation of the form:
+$\frac{\partial u}{\partial t}+\alpha u\frac{\partial u}{\partial x}+\beta\frac{\partial^{2}u}{\partial x^{2}}+\gamma\frac{\partial^{4}u}{\partial x^{4}}=0,$ (19)
+subject to periodic boundary conditions and an initial condition
+$u(0,x)=u_{0}(x).$ (20)
+The parameters $\alpha$, $\beta$, and $\gamma$ control the dynamical behavior of the equation. We use the configurations: $\alpha=100/16$, $\beta=100/162$, and $\gamma=100/164$ for chaotic behaviors.
+Appendix C Experimental Design
 We use mini-batch gradient descent, where collocation points are randomly sampled during training iteration, for all Allen-Cahn experiments, and we use full-batch gradient descent (but varying training set size) for all other experiments.
+Each full-batch collocation point sets were generated using the efficient Poisson sampling technique from Shankar (2017).
 We use the Adam optimizer Kingma & Ba (2015), and follow the learning rate schedule of Wang et al.
@@
 (2023), we also employ a learning rate annealing algorithm Wang et al.
-(2021a; 2023) to balance losses and causal training Wang et al.
+(2023) to balance losses and causal training Wang et al.
 (2022a; 2023) to mitigate causality violation in solving time-dependent PDEs and apply exact periodic boundary conditions Dong & Ni (2021) when applicable.
@@
 Random weight factorization $\mu=1.0,\sigma=0.1$
-Learning rate schedule
 Initial learning rate $10^{-3}$
@@
 Warmup steps $5\times 10^{3}$
-Training
 Training steps $3\times 10^{5}$
-Weighting
 Weighting scheme Gradient Norm Wang et al. (2022b; 2023)